OpenStudy (anonymous):

Simplify sin θ csc θ − sin2 θ.

OpenStudy (anonymous):

hint:\[\csc \theta = \frac{ 1 }{ \sin \theta }\]

OpenStudy (anonymous):

so long as \[\theta \ne n \pi, n \in \mathbb{Z}\]

OpenStudy (anonymous):

hello...

OpenStudy (anonymous):

thanks that worked

OpenStudy (anonymous):

what did you get?